I need help soving: 5+9/x=2
Found 2 solutions by user_dude2008, stanbon:
Answer by user_dude2008(1862)

5+9/x=2

9/x=-3

9=-3x

x=-3

Answer by stanbon(75887)

I need help solving: 5+9/x=2
-------
Multiply thru by x to get:
5x + 9 = 2x
3x = -9
x = -3
===============
Cheers,
Stan H.